OpenStudy (anonymous):

Use a system of equations to solve the following problem: Charles High School has a total of 94 boys and girls who play sports. If the number of boys is 22 more than twice the number of girls, how many boys play sports at this high school? 70 48 24 80

OpenStudy (anonymous):

let the number of boys be x no of girls = y x + y = 94 x = 22 + 2y now Solve both equation and find x ..

OpenStudy (angela210793):

let x be the number of boys and y the number of girls a total of 94 boys and girls can be written as: x+y=94 the number of boys is 22 more than twice the number of girls can be written as x=22 2y now